A 5kg. object travelling at 10m/s collides head-on with a 10kg. object travelling at 15m/s. If they were locked together after the collision, what is their velocity, and in what direction they will be heading?


1
Expert's answer
2021-04-12T07:00:28-0400
"MV-mv=(m+M)u\\(10)(15)-(5)(10)=(5+10)u\\u=6.67\frac{m}{s}"

They will be heading in the direction of motion of a 10kg. object.